# On the Completeness of Invariant Geometric Deep Learning Models

### v2

@@ Abstract
Invariant models, one important class of geometric deep learning models, are capable of generating meaningful geometric representations by leveraging informative geometric features in point clouds. These models are characterized by their simplicity, good experimental results and computational efficiency. However, their theoretical expressive power still remains unclear, restricting a deeper understanding of the potential of such models. In this work, we concentrate on characterizing the theoretical expressiveness of a wide range of invariant models. We first rigorously bound the expressiveness of the most classic invariant model, message-passing neural networks incorporating distance (DisGNN), restricting its unidentifiable cases to be only highly symmetric point clouds. We then show that GeoNGNN, the geometric counterpart of one of the simplest subgraph graph neural networks (subgraph GNNs), can effectively break these corner cases’ symmetry and thus achieve E(3)-completeness. By leveraging GeoNGNN as a theoretical tool, we further prove that: 1) most subgraph GNNs developed in traditional graph learning can be seamlessly extended to geometric scenarios with E(3)-completeness; 2) DimeNet, GemNet and SphereNet, three well-established invariant models, are also all capable of achieving E(3)-completeness. Our theoretical results fill the gap in the theoretical power of invariant models, contributing to a rigorous and comprehensive understanding of their capabilities. We also empirically evaluated GeoNGNN, the simplest model within the large E(3)-complete family we established, which achieves competitive results to models relying on high-order invariant/equivariant representations on molecule-relevant tasks.
@@ 1 Introduction
Learning geometric structural information from 3D point clouds constitutes a fundamental requirement for various real-world applications, including molecular property prediction, physical simulation, and point cloud classification/segmentation [51, 50, 30, 25].
@@
However, these restrictions on symmetry can impose limitations on the ability of models to approximate a broad range of symmetric functions. For example, Message Passing Neural Networks (MPNNs) [24], a representative class of Graph Neural Networks (GNNs), learn permutation symmetric functions over graphs. However, it has been shown that such functions are not universal [60, 42], leading to the development of more expressive GNN frameworks [40, 42, 62]. Similar challenges arise in the geometric setting, where models must additionally respect the Euclidean symmetry, and thus achieving geometric universality (E(3)-completeness) becomes a non-trivial task as well.
Regarding this, previous studies [58, 59, 35] have proposed complete invariants and continuous metrics for their comparison. However, these constructions do not directly correspond to neural function forms, thereby restricting their applications in downstream tasks. In neural-network context, some works [12, 48, 56, 57] proved their models can achieve geometric completeness by excluding some symmetric, difficult-to-deal cases (such as assuming non-degenerate frames). However, these cases are known to significantly impact downstream continuous function learning [46, 26]. A series of recent works [37, 10, 26, 27] proposed geometric models based on 2FWL methods [7], which initializes and updates node-pair representations iteratively, and proved their completeness. However, these models significantly diverge from the node-based MPNN paradigm, and whether a wide range of invariant models, most with weaker aggregation schemes, are also complete remains unclear. For more discussion of related works, we refer the readers to Appendix A.
@@
In this work, we systematically investigate the expressiveness of a wide range of invariant models. We begin by revisiting DisGNN, the simplest conceptual invariant model augmenting MPNNs with distance as an edge feature (SchNet [53] is a concrete implementation of it), with a specific emphasis on addressing the question: when is DisGNN powerful enough? We establish that DisGNN is a nearly-E(3)-complete model by restricting all its unidentifiable cases into a 0-measure subspace which only contains well-defined highly-symmetric point clouds. Our results extend the findings of previous works that solely concentrated on individual hand-crafted counterexamples illustrating the incompleteness of DisGNN [37, 46, 47], and study that provided a strictly coarser characterization of DisGNN’s expressiveness [26].
@@
We then show that GeoNGNN, the geometric counterpart of NGNN [63] (one simplest subgraph GNN), can effectively break all these unidentifiable cases’ symmetry through node marking, thereby achieving E(3)-completeness. Leveraging GeoNGNN as a theoretical tool, we proceed to establish the E(3)-completeness of a wide range of invariant models by showcasing their ability to implement [17] GeoNGNN. Specifically, we systematically define the geometric counterparts of all subgraph GNNs within Zhang et al. [62]’s framework by generalizing their local operations to incorporate geometric information. We then prove that all of these subgraph GNNs are E(3)-complete. Furthermore, we establish that three well-established invariant geometric models, DimeNet [20], SphereNet [39] and GemNet [22], are also all capable of achieving E(3)-completeness.
@@
These established E(3)-complete models, except for GemNet [22], are either known to be strictly weaker than 2FWL algorithm [7] in traditional graph learning or utilize less informative aggregation schemes than 2FWL-like geometric models [37, 10, 26, 27]. Nevertheless, our investigation reveals that in geometric settings, these models are equally E(3)-complete. These surprising findings can deepen our understanding of the power of invariant models, while also suggesting that the bottleneck of invariant models may lie in generalization instead of expressiveness. As a side effect, by generalizing subgraph GNNs to the geometric setting, we greatly enlarge the design space of geometric deep learning models, which we hope could inspire future, more powerful model designs.
@@ 2.1 Notations and definitions
Two point clouds are isomorphic if they are equivalent under permutation of nodes and Euclidean isometry (including rotation, translation and reflection). Invariant models explored in this work are permutation- and E(3)-invariant, meaning that they embed isomorphic point clouds into the same representation $s\in\mathbb{R}^{k}$.
We say that a model can distinguish a pair of non-isomorphic point clouds if their corresponding output $s_{1}$ and $s_{2}$ are different, and say that a model is (E(3)-)complete if it can distinguish all pairs of finite-size non-isomorphic point clouds. Since distinguishing point clouds with different node numbers is trivial, we only consider distinguishing abilities for clouds equally with finite $n$ nodes.
Instead of researching models’ ability to distinguish pairs of point clouds, we take a step further for their ability to identify each single point cloud: A model can identify a point cloud $P\in\mathbb{R}^{n\times 3}$, if for an arbitrary point cloud $P^{\prime}\in\mathbb{R}^{n\times 3}$ that is non-isomorphic to $P$, the model can always distinguish them. "Identify" allows us to understand the expressiveness of models in a finer-grained manner, since by definition, each "identify" results in infinite pairs of "distinguish". It is also obvious to see that identifying all point clouds implies E(3)-completeness.
Finally, we note that all the invariant models are parametric. However, Hornik et al. [28], Dym & Gortler [13], Amir et al. [2] showed that there exist neural function forms and parameters ensuring the injectiveness of all multiset functions within polynomial complexity, leading to lossless encoding. Under such injective constructions, neural networks can always simulate their corresponding color-refinement algorithms [60, 62, 37] which represents their maximal expressiveness. We thus directly discuss expressiveness in neural-network contexts, which implicitly assumes the injectiveness of the models. Moreover, since models are highly parametric, the completeness directly implies their universal approximation ability for all E(3)- and permutation-invariant functions [26], giving them great theoretical significance in practical applications.
@@ 2.2 DisGNN
To leverage the rich geometric information present in point clouds, a straightforward approach is to model the point cloud as a distance graph, where edges are present between nodes if their Euclidean distance is below a certain cutoff $r_{\text{cutoff}}$, and edge weights are Euclidean distance $d_{ij}$. Then, an MPNN is applied to it. We call models with such framework as DisGNN, and SchNet [53] is a representative of such works.
To be specific, DisGNN models first embed nodes based on their initial node features $X\in\mathbb{R}^{n\times d}$ (if there exists), such as atomic numbers, obtaining $h_{i}^{(0)}$ for node $i$. They then perform message passing according to the following equation iteratively:
@@
where $h_{i}^{(l)}$ represents the node embedding for node $i$ at layer $l$, $\mathcal{N}(i)$ denotes the neighbor set of node $i$. The global graph embedding is then obtained by aggregating the representations of all nodes using a permutation-invariant function $f_{\text{out}}(\{\!\!\{h_{i}^{L}\mid i\in[n]\}\!\!\})$, where $L$ is the total number of layers and $n$ is the total number of nodes.
In the theoretical analysis for DisGNN (Section 3), we assume fully-connected distance graph modeling for point clouds, i.e., $\mathcal{N}(i)$ is $[n]$ by default, and sufficient iterations until convergence. This ensures maximal geometric expressiveness of DisGNN and aligns with existing works [46, 26, 37].
@@ 3 How powerful is DisGNN?
In Pozdnyakov & Ceriotti [46], Li et al. [37], researchers have carefully constructed pairs of symmetric point clouds that DisGNN cannot distinguish to illustrate its E(3)-incompleteness. However, rather than relying on hand-crafted finite pairs of counterexamples and intuitive illustration, a more significant question arises: what properties do all the point clouds that DisGNN cannot identify share in common? Only after theoretically characterizing these corner cases can we describe and bound the expressiveness of DisGNN in a rigorous way, and then design more powerful geometric models accordingly by “solving” these corner cases.
@@
Naively, to determine whether a point cloud $P$ can be identified by DisGNN requires iterating through all other point clouds $P^{\prime}$ within the entire cloud space $\mathbb{R}^{n\times 3}$ according to its definition. Here, we propose a simple sufficient condition for a point cloud to be identifiable by DisGNN, which focuses solely on $P$. To achieve this, we first rigorously define the concept of $\mathcal{A}$-symmetry.
Definition 3.1.
@@
($\mathcal{A}$-symmetry) Given a point cloud $P\in\mathbb{R}^{n\times 3}$, let $\mathcal{A}$ be a permutation-equivariant and E(3)-invariant algorithm taking $P$ as input and produces node-level features $X^{\mathcal{A}}=\mathcal{A}(P)\in\mathbb{R}^{n\times d}$ (we use $x^{\mathcal{A}}_{i}\in\mathbb{R}^{d}$ to denote the $i$-th nodes’ feature). We define $\mathcal{A}$ center set of $P$ as $\mathcal{A}^{\text{set}}(P)=\{\frac{\sum_{i\in[n]}m(x^{\mathcal{A}}_{i})p_{i}}{\sum_{i\in[n]}m(x^{\mathcal{A}}_{i})}\mid m:\mathbb{R}^{d}\to\mathbb{R}\}$, where $m$ represents a "mass" function that maps different features to different "masses" [^2] Notice that the geometric center of $P$ is always present in $\mathcal{A}^{\text{set}}(P)$ by setting $m$ as a constant function.. Hence, $\mathcal{A}^{\text{set}}(P)$ represents the collection of "barycenters" under different masses. If $|\mathcal{A}^{\text{set}}(P)|=1$, then we say $P$ is $\mathcal{A}$-symmetric; otherwise, we say $P$ is $\mathcal{A}$-unsymmetric.
For example, we can consider $\mathcal{A}$ as the center distance encoding function $\mathcal{C}$, which computes the distance from each node to the geometric center as the node feature. A $\mathcal{C}$-unsymmetric case is shown in Figure 1(a).
@@
Figure 1: (a) A $\mathcal{C}$-unsymmetric point cloud. The blue small node represents the geometric center, and each green node represents an element in $\mathcal{C}^{\text{set}}(P)$. The large nodes constitute the point cloud, with different node colors denoting different node features. (b) Illustration of the relations among different subspaces of the point cloud. $\mathcal{A}$-sym parts encompass all $\mathcal{A}$-symmetric graphs. Each subspace has a measure of 00 and strictly contains their sub-parts (proper subset relation).
@@
Obviously, the cardinality of $\mathcal{A}^{\text{set}}(P)$ is solely dependent on the partition of node features $\mathcal{A}(P)$ and is independent of the specific feature values. Another algorithm, $\mathcal{A}$, could be $\mathcal{D}$, which applies DisGNN on point clouds until the partition of node features stabilizes. Notably, earlier studies by Delle Rose et al. [10], Li et al. [37] have demonstrated that $\mathcal{D}$ ensures a node feature partition that is no coarser compared to $\mathcal{C}$. As a consequence, $\mathcal{D}$-symmetry always implies $\mathcal{C}$-symmetry. However, determining $\mathcal{D}$-symmetry of a point cloud will need higher costs. We now present the primary theorem demonstrating that all such unsymmetric point clouds can be identified by DisGNN.
Theorem 3.2.
@@
(Unsymmetric point clouds are identifiable) Let $\mathcal{A}$ be either center distance encoding $\mathcal{C}$ or DisGNN encoding $\mathcal{D}$. Then, given an arbitrary point cloud $P\in\mathbb{R}^{n\times 3}$, $P$ is not $\mathcal{A}$-symmetric $\Rightarrow$ $P$ can be identified by DisGNN.
Key idea
@@
Given a point cloud $P$, when it has two global-level distinct anchors, such as geometric center or other centers that can be determined in a permutation-invariant and E(3)-equivariant way, DisGNN has the potential to learn triangular distance encoding for each node $i$ w.r.t. $i$ and these two global anchors, and produces complete representation for $P$ by leveraging such encoding information (Lemma E.6). For $\mathcal{C}$- or $\mathcal{D}$-unsymmetric point clouds, their center sets $\mathcal{C}^{\text{set}}(P)$ or $\mathcal{D}^{\text{set}}(P)$ have more than 1 element, within which any pair of distinct centers can be utilized by DisGNN for such encoding (Lemma E.4), thereby facilitating DisGNN to fully represent and identify $P$.
@@
We provide rigorous proof in Appendix E.2 and highly encourage readers to refer to it, where we show the superior global geometric learning ability of DisGNN and can inspire the development of new complete model designs in the triangular distance encoding perspective, as exemplified by GeoNGNN proposed in the following section.
Notice that while in Theorem 3.2 DisGNN is also applied to determine $\mathcal{D}$-symmetry, the theorem indicates that to determine the identifiability of $P$, a single execution of DisGNN on $P$ suffices, sparing the need to iterate through all possible $P^{\prime}\in\mathbb{R}^{n\times 3}$. The theorem significantly constrains the unidentifiable space of DisGNN to a notably restricted subspace, demonstrating that, though with extremely simple and naive design, DisGNN is nearly E(3)-complete. In Figure 1(b), we depict the containment relation between these well-defined subspaces. We notice that all the relations are proper subset relations, which is proved in Proposition B.1. In Appendix B, we also show the following theoretical/experimental evidence, demonstrating how restricted these subspaces are:
  • We rigorously prove that the measure of $\mathcal{C}$- and $\mathcal{D}$-symmetric clouds, as well as the unidentifiable clouds’ measure, is $\mathbf{0}$ within the whole space $\mathbb{R}^{n\times 3}$ (Theorem B.4).
  • Realizing that real-life point clouds are all sightly noisy, we investigate the proportion of $\mathcal{C}$- and $\mathcal{D}$-symmetric clouds in real-world datasets including QM9 and ModelNet40, with pre-defined tolerance errors (allowing them to be not exactly symmetric). We show that even with the largest tolerance, the proportion of unidentifiable cases amounts to less than $\mathbf{0.1}\%$ (Appendix B.3).
@@
Finally, we note that a recent study [26] also indicates the near-completeness of DisGNN, by showing that it can distinguish arbitrary pairs of point clouds from an unsymmetric point cloud $\mathbb{R}^{n\times 3}_{\text{distinct}}$. Nevertheless, our findings represent a significant advancement by introducing two strictly larger unsymmetric subsets and employing the concept of "identify" instead of "pair-wise distinguish". This provides a much finer characterization of the near completeness of DisGNN. Please refer to Appendix B for a detailed comparison.

### v3

@@ Abstract
Invariant models, one important class of geometric deep learning models, are capable of generating meaningful geometric representations by leveraging informative geometric features in point clouds. These models are characterized by their simplicity, good experimental results and computational efficiency. However, their theoretical expressive power still remains unclear, restricting a deeper understanding of the potential of such models. In this work, we concentrate on characterizing the theoretical expressiveness of a wide range of invariant models under fully-connected conditions. We first rigorously characterize the expressiveness of the most classic invariant model, message-passing neural networks incorporating distance (DisGNN), restricting its unidentifiable cases to be only highly symmetric point clouds. We then prove that GeoNGNN, the geometric counterpart of one of the simplest subgraph graph neural networks, can effectively break these corner cases’ symmetry and thus achieve E(3)-completeness. By leveraging GeoNGNN as a theoretical tool, we further prove that: 1) most subgraph GNNs developed in traditional graph learning can be seamlessly extended to geometric scenarios with E(3)-completeness; 2) DimeNet, GemNet and SphereNet, three well-established invariant models, are also all capable of achieving E(3)-completeness. Our theoretical results fill the gap in the expressive power of invariant models, contributing to a rigorous and comprehensive understanding of their capabilities.
@@ 1 Introduction
Learning geometric structural information from 3D point clouds constitutes a fundamental requirement for various real-world applications, including molecular property prediction, physical simulation, and point cloud classification/segmentation (Schmitz et al., 2019; Sanchez-Gonzalez et al., 2020; Jumper et al., 2021; Guo et al., 2020).
@@
However, these restrictions on symmetry can impose limitations on the ability of models to approximate a broad range of functions. For example, Message Passing Neural Networks (MPNNs) (Gilmer et al., 2017), a representative class of Graph Neural Networks (GNNs), learn permutation symmetric functions over graphs. However, it has been shown that such functions are not universal (Xu et al., 2018a; Morris et al., 2019), leading to the development of more expressive GNN frameworks (Maron et al., 2018; Morris et al., 2019; Zhang et al., 2023). Similar challenges arise in the geometric setting, where models must additionally respect the Euclidean symmetry, and thus achieving geometric universality (E(3)-completeness) becomes a non-trivial task as well.
@@
In this work, we systematically investigate the expressiveness of a wide range of invariant models under fully-connected conditions , where interactions occur among all points. We begin by revisiting DisGNN (Li et al., 2024), the simplest invariant model augmenting MPNNs with Euclidean distances between nodes as additional edge features. We seek to answer the question: how close is DisGNN to completeness? Through theoretical analysis, we show that DisGNN is nearly-E(3)-complete, whose unidentifiable cases can be restricted to a 0-measure subset containing well-defined highly-symmetric point clouds only. Our results extend the findings of previous works that solely concentrated on individual hand-crafted counterexamples illustrating the incompleteness of DisGNN (Li et al., 2024; Pozdnyakov & Ceriotti, 2022; Pozdnyakov et al., 2020), and a contemporary work that provided a coarser characterization of DisGNN’s expressiveness (Hordan et al., 2024b).
@@
We then show that GeoNGNN, the geometric counterpart of NGNN (Zhang & Li, 2021) (one simple subgraph GNN), can effectively break all DisGNN’s unidentifiable cases’ symmetry through node marking, thereby achieving E(3)-completeness. We then leverage GeoNGNN as a theoretical tool, proceeding to establish the E(3)-completeness of a wide range of invariant models by showcasing their ability to implement GeoNGNN. Specifically, we systematically define the geometric counterparts of subgraph GNNs within Zhang et al. (2023)’s framework by generalizing their local operations to incorporate geometric information, and then prove that all of these geometric subgraph GNNs are E(3)-complete. Furthermore, we establish that three well-established invariant geometric models, DimeNet (Gasteiger et al., 2019), SphereNet (Liu et al., 2021) and GemNet (Gasteiger et al., 2021), are also all capable of achieving E(3)-completeness.
@@
These established E(3)-complete models, except for GemNet (Gasteiger et al., 2021), are either known to be strictly weaker than 2-FWL (Cai et al., 1992) in traditional graph learning setting, or utilize less informative aggregation schemes than 2-FWL-like geometric models (Li et al., 2024; Delle Rose et al., 2023; Hordan et al., 2024b; a) (See Section 5 for details). However, our investigation reveals that in the geometric setting, these models are equally E(3)-complete. These surprising findings can deepen our understanding of the power of invariant models, while also suggesting that the bottleneck of invariant models may lie in generalization instead of expressiveness. As a side effect, by generalizing subgraph GNNs to the geometric setting, we greatly enlarge the design space of geometric deep learning models, which we hope could inspire future, more powerful model designs.
2 Related Works
Designing complete invariant descriptors for point clouds plays a crucial role for scientific problems (Nigam et al., 2024), since complete invariants directly allow universal function approximations when coupled with MLPs (Hordan et al., 2024b; Li et al., 2024). Previous works, such as Widdowson & Kurlin (2022; 2023); Kurlin (2023), have proposed polynomial-time algorithms to compute and compare complete invariants that uniquely determine point clouds. However, these invariants are often highly structured, such as sets of matrices (Kurlin, 2023), and the absence of corresponding neural-network formulations limits their practical applicability.
In the neural-network context, most prior works propose powerful models but typically with “weaker completeness”(Wang et al., 2022; Puny et al., 2021; Duval et al., 2023; Du et al., 2024; Villar et al., 2021). For example, ComENet (Wang et al., 2022) uses nearest neighbors as references for computing higher-order geometric information, which is infeasible for symmetric point clouds where nodes can have multiple nearest neighbors, thereby limiting its completeness over a asymmetric subset of point clouds  [^1] Arbitrarily selecting one when multiple nearest neighbors exist could break permutation invariance, a fundamental principle ensuring stable outputs. We focus only on invariant models.. Frame-based methods, such as FAENet (Duval et al., 2023), LEFTNet (Du et al., 2024), and others (Wang & Zhang, 2022; Puny et al., 2021), project coordinates or vector features onto local or global equivariant frames and employ powerful structures like MLPs (Hornik et al., 1989) to ensure expressiveness. However, these frames can degenerate in symmetric structures (e.g., structures with non-distinct eigenvectors (Puny et al., 2021; Duval et al., 2023)), a limitation that these methods do not carefully address and theoretically characterize. Although symmetric point clouds often represent zero-measure cases (Hordan et al., 2024b), they can significantly impact downstream continuous function learning, as demonstrated in (Pozdnyakov & Ceriotti, 2022; Hordan et al., 2024b).
Several previous works (Dym & Maron, 2020; Joshi et al., 2023; Hordan et al., 2024b; Li et al., 2024; Delle Rose et al., 2023; Hordan et al., 2024a) have characterized geometric models’ expressiveness in a more rigorous sense. The seminal work (Dym & Maron, 2020) proved the universality of structures like TFN (Thomas et al., 2018), while such universality requires arbitrarily high-order tensors that are computationally expensive in practice. Another related work, GWL (Joshi et al., 2023), offers many insightful theoretical conclusions, including that GWL provides upper bounds on the expressiveness of many equivariant models (Batatia et al., 2022; Satorras et al., 2021). However, GWL relies on injective multiset functions involving equivariant features, which do not exhibit practical polynomial-time neural forms like scalar multiset functions proposed in (Dym & Gortler, 2024; Amir et al., 2024), thus serving more as conceptual upper bounds. Joshi et al. (2023) also does not directly address completeness of models as we do. A recent work (Cen et al., 2024) characterizes the symmetry of point clouds from point group perspective, and investigate the expressiveness through degeneration of equivariant functions. Notably, recent works (Li et al., 2024; Delle Rose et al., 2023; Hordan et al., 2024b; a) have proposed provably complete geometric models that can produce complete invariant features. Nevertheless, these methods are predominantly built upon the 2-FWL framework (Cai et al., 1992), and the completeness of a broader class of invariant models—many of which employ weaker aggregation schemes (Gasteiger et al., 2019)—remains an open question. In this work, we focus on a significantly broader range of invariant methods (Schütt et al., 2018; Gasteiger et al., 2019; 2021; Liu et al., 2021; Zhang & Li, 2021; Zhang et al., 2023), some of which are newly proposed based on subgraph GNNs in traditional graph learning (Zhang & Li, 2021; Zhang et al., 2023), with the aim to fully characterize their completeness rigorously.
@@ 3.1 Notations and Definitions
Two point clouds $P_{1}\in\mathbb{R}^{n_{1}\times 3}$ and $P_{2}\in\mathbb{R}^{n_{2}\times 3}$ are isomorphic if $n_{1}=n_{2}=n$, and there exists a rotation matrix $\mathbf{R}\in O(3)$, a translation vector $\mathbf{t}\in\mathbb{R}^{3}$, and a permutation matrix $\mathbf{P}\in\mathbb{R}^{n\times n}$ such that $(\mathbf{P}P_{1})\mathbf{R}+\mathbf{t}=P_{2}$. Invariant models considered in this study are all permutation- and E(3)-invariant, meaning they embed isomorphic point clouds into the same $k$-dimensional representation $s\in\mathbb{R}^{k}$.
Definition 3.1 (Distinguish, Identify, and E(3)-Completeness).
Let $f:\bigcup_{n=1}^{\infty}\mathbb{R}^{n\times 3}\to\mathbb{R}^{k}$ be a permutation- and E(3)-invariant function that maps point clouds to a $k$-dimensional representation. We define the following concepts:
  • Distinguish: Given two non-isomorphic point clouds $P_{1}$ and $P_{2}$, if $f(P_{1})\neq f(P_{2})$, we say $f$ can distinguish $P_{1}$ and $P_{2}$.
  • Identify: For a point cloud $P_{1}$, if for any non-isomorphic $P_{2}$, we always have $f(P_{1})\neq f(P_{2})$, we say $f$ can identify $P_{1}$.
  • E(3)-Completeness: If for any pair of non-isomorphic point clouds $P_{1}$ and $P_{2}$, we always have $f(P_{1})\neq f(P_{2})$, we say $f$ is E(3)-complete.
Notably, we introduce the novel concept of “Identify”, which allows for a finer-grained analysis of model expressiveness than commonly adopted “Distinguish”: By definition, each “identify” operation encompasses an infinite number of “distinguish” pairs. It is evident that identifying all point clouds implies E(3)-completeness. Furthermore, since distinguishing point clouds with different node counts is straightforward, we focus solely on the case where point clouds have the same finite size.
For a parametric model $f_{\theta}$, we investigate its maximal expressive form where all intermediate functions are injective, following typical ways (Delle Rose et al., 2023; Li et al., 2024). Notably, recent work by Dym & Gortler (2024); Amir et al. (2024) proves the existence of neural function forms and parametrizations that ensure injectivity for multiset functions commonly used in geometric models (Gasteiger et al., 2019; 2021; Li et al., 2024), allowing us to achieve such completeness in polynomial time and under (for Lebesgue) almost every parametrization.
@@ 3.2 DisGNN
To leverage the rich geometric information present in point clouds, a straightforward approach is to model the point cloud as a distance graph, where edges are present between nodes if their Euclidean distance is below a certain cutoff $r_{\text{cutoff}}$, and edge weights are Euclidean distance $d_{ij}$. Then, an MPNN is applied to it. We call models with such framework as DisGNN, and SchNet (Schütt et al., 2018) is a representative of such works.
To be specific, DisGNN models first embed nodes based on their initial node features $X\in\mathbb{R}^{n\times d}$ (if there exists), such as atomic numbers, obtaining initial node embedding $h_{i}^{(0)}$ for node $i$. They then perform message passing according to the following equation iteratively:
@@
where $h_{i}^{(l)}$ represents the node embedding for node $i$ at layer $l$, $\mathcal{N}(i)$ denotes the neighbor set of node $i$. The global graph embedding is then obtained by aggregating the representations of all nodes using a permutation-invariant function $f_{\text{out}}(\{\!\!\{h_{i}^{(L)}\mid i\in[n]\}\!\!\})$, where $L$ is the total number of layers and $n$ is the total number of nodes.
In the theoretical analysis for DisGNN (Section 4), we assume fully-connected distance graph modeling for point clouds, i.e., $\mathcal{N}(i)$ is $[n]$ by default, and sufficient iterations until convergence. This ensures maximal geometric expressiveness of DisGNN and aligns with existing works (Pozdnyakov & Ceriotti, 2022; Hordan et al., 2024b; Li et al., 2024).
@@ 4 How Powerful is DisGNN?
In Pozdnyakov & Ceriotti (2022); Li et al. (2024), researchers have carefully constructed pairs of symmetric point clouds that DisGNN cannot distinguish to illustrate its E(3)-incompleteness. However, rather than relying on hand-crafted finite pairs of counterexamples and intuitive illustration, a more significant question arises: what properties do the point clouds that DisGNN cannot identify share in common? Only after theoretically characterizing these corner cases can we describe and bound the expressiveness of DisGNN in a rigorous way, and then design more powerful geometric models accordingly by “solving” these corner cases.
@@
Naively, to determine whether a point cloud $P$ can be identified by DisGNN requires iterating through all other point clouds $P^{\prime}$ within the entire cloud set $\mathbb{R}^{n\times 3}$ according to its definition. Here, we propose a simple sufficient condition for a point cloud to be identifiable by DisGNN, which focuses solely on $P$. To achieve this, we first rigorously define the concept of $\mathcal{A}$-symmetry.
Definition 4.1 ($\mathcal{A}$-symmetry).
@@
Given a point cloud $P\in\mathbb{R}^{n\times 3}$, let $\mathcal{A}$ be a permutation-equivariant and E(3)-invariant algorithm taking $P$ as input and produces node-level features $X^{\mathcal{A}}=\mathcal{A}(P)\in\mathbb{R}^{n\times d}$ (we use $x^{\mathcal{A}}_{i}\in\mathbb{R}^{d}$ to denote the $i$-th nodes’ feature). We define $\mathcal{A}$ center set of $P$ as $\mathcal{A}^{\text{set}}(P)=\{\frac{\sum_{i\in[n]}m(x^{\mathcal{A}}_{i})p_{i}}{\sum_{i\in[n]}m(x^{\mathcal{A}}_{i})}\mid m:\mathbb{R}^{d}\to\mathbb{R},\sum_{i\in[n]}m(x^{\mathcal{A}}_{i})\neq 0\}$. If $|\mathcal{A}^{\text{set}}(P)|=1$, then we say $P$ is $\mathcal{A}$-symmetric; otherwise, we say $P$ is $\mathcal{A}$-asymmetric.
Intuitively, $\mathcal{A}$ assigns features to nodes of the point cloud based on their spatial properties, while $m$ represents a “mass” function that maps these features to corresponding “masses”. [^3] Note that the geometric center of $P$ is always included in $\mathcal{A}^{\text{set}}(P)$ by defining $m$ as a constant function. Thus, $\mathcal{A}^{\text{set}}(P)$ denotes the collection of “barycenters” associated with different “mass” assignments. For example, we can consider $\mathcal{A}$ as the distance encoding function, denoted as $\mathcal{C}$, which computes the distance from each node to the geometric center as the node feature. A $\mathcal{C}$-asymmetric case is shown in Figure 1(a).
@@
Figure 1: (a) A $\mathcal{C}$-asymmetric point cloud. The blue small node represents the geometric center, and each green node represents an element in $\mathcal{C}^{\text{set}}(P)$. The large nodes constitute the point cloud, with different node colors denoting different node features. (b) Illustration of the relations among different subsets of the point cloud. $\mathcal{A}$-sym parts encompass all $\mathcal{A}$-symmetric graphs. Each subset has a measure of 00 and strictly contains their sub-parts (proper subset relation).
@@
Obviously, the cardinality of $\mathcal{A}^{\text{set}}(P)$ is solely dependent on the partition of node features $\mathcal{A}(P)$ and is independent of the specific feature values. We also propose a more powerful algorithm, DisGNN encoding, denoted as $\mathcal{D}$, which applies DisGNN to point clouds iteratively until the partition of node features stabilizes. Notably, earlier studies by Delle Rose et al. (2023); Li et al. (2024) have shown that $\mathcal{D}$ guarantees a node feature partition that is no coarser than that of $\mathcal{C}$. Consequently, $\mathcal{D}$-symmetry always implies $\mathcal{C}$-symmetry, representing a stricter form of symmetry.
Based on these definitions, we now present the primary theorem demonstrating that all such asymmetric point clouds can be identified by DisGNN.
Theorem 4.2 (Asymmetric point clouds are identifiable).
@@
Let $\mathcal{C}$ denote the center distance encoding and $\mathcal{D}$ the DisGNN encoding. Then, given an arbitrary point cloud $P\in\mathbb{R}^{n\times 3}$, $P$ is $\mathcal{C}$-asymmetric $\Rightarrow$ $P$ is $\mathcal{D}$-asymmetric $\Rightarrow$ $P$ can be identified by DisGNN.
@@
Key idea. Given a point cloud $P$, when it has two global-level distinct anchors, such as geometric center or other centers that can be determined in a permutation-invariant and E(3)-equivariant way, DisGNN has the potential to learn triangular distance encoding for each node $i$ w.r.t. $i$ and these two global anchors, and produces complete representation for $P$ by leveraging such encoding information (Lemma B.6). For $\mathcal{C}$- or $\mathcal{D}$-asymmetric point clouds, their center sets $\mathcal{C}^{\text{set}}(P)$ or $\mathcal{D}^{\text{set}}(P)$ have more than 1 element, within which any pair of distinct centers can be utilized by DisGNN for such encoding (Lemma B.4), thereby facilitating DisGNN to fully represent and identify $P$.
@@
We provide rigorous proof in Appendix B.3 and encourage readers to refer to it, where we show the superior global geometric learning ability of DisGNN and can inspire the development of new complete model designs in the triangular distance encoding perspective, as exemplified by GeoNGNN proposed in the following section.
Theorem 4.2 shows that to test if a point cloud is identifiable by DisGNN, one can first check its $\mathcal{C}$-symmetry, which is quick and intuitive. If it is $\mathcal{C}$-asymmetric, it is identifiable; otherwise, one can test for the more powerful but costlier $\mathcal{D}$-symmetry. Notably, although the latter test still requires running DisGNN, it demonstrates that a single execution of DisGNN on $P$ can provide an indication of identifiability, offering insights without necessitating an exhaustive search over all $P\textquoteright\in\mathbb{R}^{n\times 3}$.
Importantly, Theorem 4.2 restricts the unidentifiable set of DisGNN to a constrained, highly symmetric subset, as illustrated in Figure 1(b). In Theorem 4.3, we provide a rigorous analysis demonstrating that the unidentifiable set of DisGNN has zero measure. These results sufficiently demonstrate that, despite its extremely simple and naive design, DisGNN is nearly E(3)-complete.
Theorem 4.3 (Unidentifiable set of DisGNN has measure zero).
The Lebesgue measure on $\mathbb{R}^{n\times 3}$ of the $\mathcal{C}$-symmetric, $\mathcal{D}$-symmetric, and unidentifiable point cloud sets is zero.
@@
Finally, we note that a recent study (Hordan et al., 2024b) also indicates the near-completeness of DisGNN, by showing that it can distinguish arbitrary pairs of point clouds from an asymmetric point cloud $\mathbb{R}^{n\times 3}_{\text{distinct}}$. Nevertheless, our findings represent a significant advancement by introducing a strictly larger asymmetric subset and employing the concept of “identify” instead of “pair-wise distinguish”. This provides a finer characterization of the near completeness of DisGNN. Please refer to Section C.2 for a detailed comparison.